Perdon Corporation manufactures safes-large mobile safes, and large walk-in stationary bank safes. As part of its annual budgeting process, Perdon is analyzing the proftability of its two products. Part of this analysis involves estimating the amount of overhead to be allocated to each product line. The information shown below relates to overhead Mobile Safes 190 300 Walk-in Safes 54 200 Units planned for production Material moves per product line Purchase orders per product line Direct labor hours per product line 20340 790 1,690 x Your answer is incorrect. The total estimated manufacturing overhead was $ 272,000. Under traditional costing (which assigns overhead on the basis of direct labor hours) what amount of manufacturing overhead costs are assigned to: (Round answers to 2 decimal places, eg, 12.25) (1) One mobile safe $ (2) One walk-in safe $ 143158 per unit 503704 per unit
Answer #1

Total labor hours = 790+1690 = 2480 Hours

Calculate overhead assigned :

1) One mobile safe = 272000*790/2480 = $86645/190 = $456.03 per unit

2) One Walk in safe = 272000/1690/2480 = $185355/54 = $3432.50 per unit